# Free-Lunch Color-Texture Disentanglement for Stylized Image Generation

### v2

@@ 3.2 Style Attributes Disentanglement
Here, we present our Style Attributes Disentanglement method (SADis). An overview is provided in Fig. 3. To successfully extract the color and texture information, we exploit the image-prompt additivity property, explained in Sec. 3.2.1. Next, we explain the color-texture extraction (CTE) in Sec. 3.2.2, and the noise-regularized whitening-coloring transformation (RegWCT) is detailed in Sec. 3.2.3.
@@ 3.2.1 Image-Prompt Additivity
To achieve disentangled stylization, we first analyze the additivity property within image prompts, which can be derived using the CLIP image encoder, as also adopted in the IP-Adapter [75]. This property, which has been explored in text prompt spaces [51, 6], also holds in the image prompt space. As an example in Fig. 4-(Left), we subtract the grayscale image embedding from the original color image embedding, that results in the T2I generation with similar color palettes. We show additional analysis with subtracting objects from the given images in the Supplementary Material. We refer to this property as Image-Prompt Additivity. Using this property, we extract color and texture information from reference images and apply these attributes effectively in T2I generation.
@@ 3.2.2 Color-Texture Extraction
where $\mathbf{GS}$ is the grayscale operation. This strips away semantic information to retain only the color attributes.
For texture extraction, we first convert the texture reference image to grayscale to remove any influence of its color palette, which is formulated as $\mathbf{Emb}_{tx}^{*}=\tau_{\phi}(\mathbf{GS}(\mathcal{I}_{tx}))$, However, using only the grayscale texture embedding creates a mismatch in scale between the color and texture branches, often resulting in overly gray tones in T2I generations (as shown in Fig. 5-2nd col.). To address this, we average the gray texture image to obtain a pure gray image, which summarizes the main gray tone information in the $\mathbf{GS}(\mathcal{I}_{tx})$ image. Following that, we concatenate both embeddings to obtain the initial texture representations:
@@
Next, we conduct the Singular-Value Decomposition (SVD) over the texture representations. Inspired by [21, 43], we assume that the main singular values of $\mathbf{Emb}_{tx}^{{}^{\prime}}$ correspond to the shared fundamental information of these two grayscale images, specifically the grayscale tone. We then have: $\mathbf{Emb}_{tx}^{{}^{\prime}}=\textbf{\emph{U}}{\boldsymbol{\Sigma}}{\textbf{\emph{V}}}^{T}$, where $\boldsymbol{\Sigma}=diag(\sigma_{0},\sigma_{1},\cdots,\sigma_{n_{j}})$, the singular values $\boldsymbol{\sigma}_{0}\geq\cdots\geq\boldsymbol{\sigma}_{n_{j}}$ To suppress the gray tone expression and extract the texture information, we introduce the augmentation for each singular value as:
@@
where $e$ is the exponential, $\gamma$ and $\beta$ are parameters with positive numbers. We then recover it as $\mathbf{Emb}_{tx}=\textbf{\emph{U}}{\hat{\boldsymbol{\Sigma}}}{\textbf{\emph{V}}}^{T}$, with the updated $\hat{\boldsymbol{\Sigma}}=diag(\hat{\sigma_{0}},\hat{\sigma_{1}},\cdots,\hat{\sigma_{n_{j}}})$. That effectively reduces residual gray-color influences while preserving texture details. After extracting both color and texture reference image embeddings, we achieve a baseline of the DisIG problem. The T2I inference process is now formulated as: $z_{t-1}^{{}^{\prime}}=\mathcal{G}(z_{t},t,\mathcal{C}_{text},\mathbf{Emb}_{tx}\oplus\mathbf{Emb}_{clr})$. Note that, we only inject the disentangled embedding $\mathbf{Emb}_{tx}\oplus\mathbf{Emb}_{clr}$ into the first decoder layer to compute cross-attention maps, which shows better stylization performance as previous works proved [67, 68, 2, 66]. Examples shown in Fig. 5 (2nd and 3rd cols.) demonstrate the effectiveness of our proposal.
Figure 5: We suppress the gray expression by introducing extra SVD reweighting to the texture representations.
@@ 3.2.3 Regularized Whitening-Coloring Transforms
WCT: whitening-coloring transformation. Solely applying our CTE method does not fully ensure a precise color palette match between the reference color image $\mathcal{I}_{clr}$ and the generated output. We hypothesize that while CLIP embeddings capture high-level semantic information, they may not fully represent nuanced color distribution details. To better align the color distributions of the generated image with the color reference, we incorporate a whitening-coloring transform (WCT) [44, 10, 31] applied to the noisy latent as $z_{t}^{wct}=\mathbf{WCT}(z_{t}^{{}^{\prime}})$. Please refer to the Supplementary for detailed WCT formulas. While traditionally used in style transfer, we are the first to apply this transformation directly within the latent space of a T2I diffusion model. This approach allows for a more faithful color transfer by aligning the latent representations with the reference color palette, enhancing the stylistic consistency of generated images. Examples of applying the WCT transform during the T2I generation process are presented in Fig. 6-(a), demonstrating an improved alignment of color palettes between the generated images and the reference color images. This enhancement visually confirms the effectiveness of WCT in achieving more accurate color transfer, resulting in a closer match to the desired color distribution.
@@
Noise Regularization for WCT transform. However, as shown in Fig. 6-(a), the WCT process tends to compress texture details, distorting high-frequency information, which is also observed in previous works [80, 42]. This issue is attributed to the signal-leak bias [16] and the inverted latent distribution gap from the inversion process [80]. These terms essentially describe the same underlying phenomenon: during training, the distribution of $z_{T}$—the latent representation after adding Gaussian noise to $z_{0}$—does not perfectly match a standard Gaussian. Instead, it always contains the leakage or some certain prior information towards the original image, causing the residual structure to persist even at high noise levels. This biased noise in the forward diffusion process leads to a slight misalignment in texture representation during generation. To address this, we propose adding a small-scale noise to the latent representation as $z_{t}=z_{t}^{wct}+\lambda\cdot\delta,\delta\sim\mathcal{N}(0,1)$, determined by a scale hyperparameter $\lambda$, to recover the lost high-frequency details. By introducing this noise regularization term, the regularized whitening-coloring transformation (RegWCT) achieves improved generative performance, better aligning both color and texture in the generated images with the reference images. Given a latent $z_{t}^{{}^{\prime}}$ at timestep $t$, the color rectified latent $z_{t}$ after RegWCT is formulated as $z_{t}=\left(1-\omega\right)z_{t}^{{}^{\prime}}+\omega\cdot\mathbf{RegWCT}(z_{t}^{{}^{\prime}})$, where $\omega$ acts as a balance weight. To prevent the inference trajectory from deviating too much from the original one, we only perform the RegWCT transformation during the intermediate steps as $[T_{start},T_{end}]$. With all these techniques, including color-texture extraction (CTE) and Regularized WCT transformation (RegWCT), our method SADis achieves customizable and flexible Disentangled Stylized Image Generation (as shown in Fig. 5-4th col.).
Figure 6: (top) We enable preciser color alignment with $\mathcal{I}_{clr}$ through RegWCT, while WCT manipulation compresses texture details. (bottom) The distribution of the image latent generated with RegWCT technique is closer to that of the image generated without WCT in the frequency domain, across diverse time steps.
@@ 3.2.4 ControlNet-based real image stylization
Furthermore, our method SADis can be integrated with ControlNet [77] $\mathcal{CN}$ to facilitate image-based stylized generation, as illustrated in Fig. 11 (Left). By utilizing any pretrained ControlNet model (e.g. Canny-conditioned) as the base pipeline while maintaining all other hyperparameters consistent with SADis, we are able to significantly broaden the applicability of our method SADis to real image cases. More specifically, we have an input $\mathbf{I}_{c}$ as the content image, it is passed through the ControlNet $\mathcal{CN}(\mathbf{I}_{c})$ as conditions for T2I generation model as $z_{t-1}^{{}^{\prime}}=\mathcal{G}(z_{t},t,\mathcal{CN}(\mathbf{I}_{c}),\mathbf{Emb}_{tx}\oplus\mathbf{Emb}_{clr})$, where $z_{T}\sim\mathcal{N}(0,1)$ and the textual prompt as null $\mathcal{P}=\text{``''}$ “”.
Figure 7: Disentangled Stylized Image Generation (DisIG) performance compared with baseline methods. For other approaches, we used GPT-4o to generate color descriptions based on the color reference image, incorporating these descriptions into the text prompts (details shown in the Supplementary). Our SADis accepts separate color and texture reference images to achieve flexible control.
@@ 4.1 Experimental Setups
Datasets. To ensure a fair comparison, we randomly select 40 images in total from the WikiArt [64] and StyleDrop [62] datasets. For our method, SADis, each of these images serves as either a color reference or a texture reference to enable color-texture disentanglement. In contrast, since the comparison methods lack disentanglement capabilities, we treat each of these 40 images as a style reference for these methods, supplemented by auxiliary text prompts generated from GPT-4o [1] as a strong captioning model. For quantitative comparisons, we used 20 images as the color reference set and 20 as the texture reference set. We also sampled 10 content prompts from StyleDrop, resulting in 4,000 stylized images per method to ensure fair and extensive comparisons with numerous images.
@@
Evaluation Metrics. We evaluate our method using multiple quantitative metrics to ensure a comprehensive assessment. (1) For a general quality evaluation, we use CLIP-Score (CLIP) [28] to evaluate the T2I generation performance, specifically measuring the semantic consistency between the generated image and the text prompt. (2) For the evaluation of color attribute alignment, the MS-SWD [25] metric specifically evaluates color distance between generated and reference images. Color histogram distance (C-Hist) computes the distance between both color histograms. The GPT-4o color score [1] requests the GPT model to rate from 0-5 according to the color consistency between the color reference and the generated image. (3) For the texture quality, the Kernel-Inception Distance (KID) [5] is assessing the quality of generated images by measuring the dissimilarity between the real and generated image distributions [^5] We use feature dimension as 2048 and evaluate on the 4000 images.. CLIP-I is used to evaluate the similarity between the texture image and the generated image in grayscale.
Figure 8: SADis remains effective when exchanging texture and color images (left) or varying color image properties such as saturation and illuminant (right). This further demonstrates the robustness of SADis in realistic DisIG scenario.
@@
Implementation Details. We build our method SADis upon the SDXL [53]. To apply the image embeddings as additional conditions to the T2I model, we use the IP-Adapter [75] pretrained projectors. Both of these models are based on the CLIP model [55], where the SDXL model utilizes the CLIP text encoder to generate textual embeddings and the IP-Adapter leverages the CLIP image encoder to extract image embeddings before the embedding projector. For the hyperparameters, we set $\lambda=0.01,T_{start}=0.8T,T_{end}=0.6T,\omega=0.5,\gamma=0.003,\beta=1.0$.
@@
Comparison Methods. We compare with several state-of-the-art tuning-free stylization approaches based on the T2I diffusion models: DEADiff [54], InstantStyle [67], IP-Adapter [75], CSGO [73], DreamStyler [3], StyleDrop [62] and Artist [39]. For these baseline methods, we utilize the GPT-4o [1] model to generate captions from the color reference image, which are appended to the original textual prompt to guide the T2I generation. Additionally, SDXL [53] and Artist model are included in comparisons by generating captions from both the color and texture reference images, appending these captions to the textual prompts to guide the generation process. We include much more experimental setup details, qualitative and quantitative results in the Supplmentary Material.
@@ 4.2 Experimental Results
Qualitative Comparison. In our comparison of generation quality, visual results under the Disentangled Stylized Image Generation (DisIG) scenario are shown in Fig. 7 and  Fig. 9. While baseline methods manage to capture some style information from the style reference image (the texture image for SADis), they fail to represent the color palette provided by the color reference accurately. The IP-Adapter and InstantStyle methods even retain original layouts and elements from the style images, deviating from the intended stylization. Other baselines, such as Artist [39] and DEADiff [54], exhibit poor generation quality, reflecting a limited understanding of both color and texture attributes. In contrast, our method, SADis, allows for both texture and color reference inputs, offering precise, controllable T2I stylization. The generated images by SADis exhibit significantly improved color fidelity and texture detail compared to other approaches, demonstrating its effectiveness in disentangled style attribute transfer. To further validate SADis’s robustness, we conduct two tests: (1) swapping the color and texture reference images, and (2) modifying their saturation and illumination levels, as shown in Fig. 8. Unlike existing approaches such as InstantStyle, which cannot preserve image content under these conditions (detailed in Supplementary), SADis successfully applies color and texture independently while keeping the content largely unchanged.
Method CLIP Color Texture
@@
SDXL [53] 0.272 9.51 1.20 3.01 0.69 0.08
IPAdapter [75] 0.233 11.54 1.23 2.84 0.84 0.043
InstantStyle [67] 0.261 12.53 1.32 2.80 0.74 0.056
Artist [39] 0.269 10.48 1.39 2.82 0.69 0.089
DEADiff [54] 0.267 11.20 1.24 2.73 0.69 0.087
StyleDrop [62] 0.275 13.52 1.43 2.67 0.70 0.054
DreamStyler [3] 0.277 12.17 1.26 2.39 0.71 0.060
CSGO [73] 0.280 14.25 1.36 2.63 0.69 0.071
SADis (Ours) 0.281 5.57 0.96 3.34 0.74 0.049
Table 1: Quantitative Comparison. The best and second-best numbers are marked with bold and underlined respectively.
Method Color Texture
@@
SADis (Ours) 5.57 0.96 0.74
$\boldsymbol{-}$ SVD 5.70 1.01 0.76
$\boldsymbol{-}$ RegWCT 8.05 1.06 0.75
$\boldsymbol{-}$ SVD $\boldsymbol{-}$ RegWCT 8.93 1.10 0.76
Table 2: Ablation by removing each components of SADis.
Figure 10: User study on comparison methods: evaluating color, texture, and combined color&textue alignment.
@@
Quantitative Comparison. A detailed quantitative comparison is provided in Tab. 1 to further support our findings. Our method, SADis, retains text-image alignment quality at a level comparable to the base SDXL model [53], as indicated by the CLIP score. This comparison demonstrates that SADis preserves alignment with the textual prompt better than other methods. Regarding color alignment, SADis significantly outperforms other approaches. For texture representation, the IP-Adapter-based methods, including InstantStyle [67] and SADis, show superior performance over other baselines. They achieve higher texture-related metrics at the expense of color fidelity and text-image alignment quality. The high texture scores for IP-Adapter can be explained by the fact that they control all cross-attention layers; However, this leads to high-semantic content leakage (like the same stars in the Van Gogh bear in Fig. 7). We decided to prevent this by only controlling the cross-attention of the low-level layers following InstantStyle [67, 2, 66]; Therefore, our texture results are close to those of InstantStyle, but at the same time we greatly improve color fidelity. In addition, their generated images follow the same structure as the original texture image, showing that they do not disentangle structure from texture but overfit to the contents of the texture images.
@@
Ablation Study. The ablation study for each component of SADis is listed in Tab. 2. Results show that SVD rescaling in the CTE process and the noise-regularized RegWCT techniques significantly enhance color alignment while only slightly reducing texture precision, which is almost undetectable in the CLIP-I metric. The trade-off between texture and color alignment is optional for users, allowing them to adjust the balance based on the requirements.
@@
User Study. To better assess alignment with human preferences, as shown in Fig. 10, we conducted a user study with 24 participants (30 sextuplets/user), collecting 720 data for each method. Each one was asked to “select the best image from pairs of generated images, taking into account overall quality (considering both color and texture alignments), color alignment, and texture alignment, respectively.” Our method, SADis, significantly outperformed other baseline approaches, with at least a 20% improvement in preference. That finally demonstrates the potency of SADis and high alignment with human preference.
4.3 Additional Applications
Image-based stylization. SADis Integrated with ControlNet achieves image-based stylized generation, as shown in Fig. 11(left). Here, we adopt ControlNet (Canny) as the base model, setting its conditioning scale to 0.6. This integration broadens the application scenarios of SADis.

### v3

@@ 3.3 Style Attributes Disentanglement
Here, we present our Style Attributes Disentanglement method (SADis). An overview is provided in Fig.˜3. To successfully extract the color and texture information, we exploit the image-prompt additivity property as in Sec. 3.3.1. Next, we explain the color-texture extraction (CTE) in Sec.˜3.3.2, and the noise-regularized whitening-coloring transformation (RegWCT) is detailed in Sec.˜3.3.3.
@@ 3.3.1 Image-Prompt Additivity
To achieve disentangled stylization, we first analyze the additivity property of image prompts. This property is the main insight on which our color-texture disentanglement is based. It is illustrated using the CLIP image encoder, as adopted in the IP-Adapter ye2023ip-adapter . Note that this property has been explored in text prompt spaces NIPS2013_wordembedding ; manuel2023sega ; here we show that it also holds in the image embedding space.
Consider the image embedding space as shown in Fig.˜4-(a). Here we show a PCA plot of the embedding space for a distribution of images generated with three different text prompts. The embedding of a color image contains all information of the image (see e.g. points $A$ and $B$). If we compare this embedding with the embedding of its grey-scale version ($A^{\prime}$, $B^{\prime}$), the difference between the two embeddings is due to the removal of color information. We here hypothesize that adding this difference to the embedding of another grey image, would transfer the color information to this image. This is illustrated for the figure of the grey bird image $C$. We can also see that the differences of $A-A^{\prime}$ and $B-B^{\prime}$ form a relatively compact distribution.
@@ 3.3.2 Color-Texture Extraction
where $\mathbf{GS}$ is the grayscale operation, $\mathbf{Emb}_{clr}\in\mathbb{R}^{n_{t}\times c}$. This strips away semantic information to retain only the color attributes (See Fig.4-b).
For texture extraction, we first convert the texture reference image to grayscale to remove any influence of its color palette, which is formulated as $\mathbf{Emb}_{tx}^{*}=\tau_{\phi}(\mathbf{GS}(\mathcal{I}_{tx}))$, $\mathbf{Emb}_{tx}^{*}\in\mathbb{R}^{n_{t}\times c}$. However, using only the grayscale texture embedding creates a mismatch in scale between the color and texture branches, often resulting in overly gray tones in T2I generations (as shown in Fig.˜5 (a)-2nd col.). To address this, we average the gray texture image to obtain a pure gray image, which summarizes the main gray tone information in the $\mathbf{GS}(\mathcal{I}_{tx})$ image. Following that, we concatenate both embeddings from token-wise to obtain the initial texture representations:
@@
$\mathbf{Emb}_{tx}^{{}^{\prime}}\in\mathbb{R}^{2n_{t}\times c}$. Next, we conduct the Singular-Value Decomposition (SVD) over the texture representations. Inspired by (gu2014weighted, ; li2023get, ), we assume that the main singular values of $\mathbf{Emb}_{tx}^{{}^{\prime}}$ correspond to the shared fundamental information of these two grayscale images, specifically the grayscale tone. We then have: $\mathbf{Emb}_{tx}^{{}^{\prime}}=\textbf{\emph{U}}{\boldsymbol{\Sigma}}{\textbf{\emph{V}}}^{T}$, where $\boldsymbol{\Sigma}=diag(\sigma_{0},\sigma_{1},\cdots,\sigma_{n_{t}},\cdots,\sigma_{2n_{t}-1})$,the singular values $\boldsymbol{\sigma}_{0}\geq\cdots\geq\boldsymbol{\sigma}_{2n_{t}-1}$. To suppress the gray tone expression and extract the texture information, we introduce the augmentation for each singular value as:
@@
where $e$ is the exponential, $\gamma$ and $\beta$ are parameters with positive numbers. We then recover it as $\mathbf{Emb}_{re}=\textbf{\emph{U}}{\hat{\boldsymbol{\Sigma}}}{\textbf{\emph{V}}}^{T}$, with the updated $\hat{\boldsymbol{\Sigma}}=diag(\hat{\sigma_{0}},\hat{\sigma_{1}},\cdots,\hat{\sigma}_{2n_{t}-1})$. The texture embedding is obtained by $\mathbf{Emb}_{tx}=\mathbf{Emb}_{re}[\,:n_{t},\,:\,]$. That effectively reduces residual gray-color influences while preserving texture details. After extracting both color and texture reference image embeddings, we achieve a baseline of the DisIG problem. The T2I inference process is now formulated as: $z_{t-1}^{{}^{\prime}}=\mathcal{G}(z_{t},t,\mathcal{C}_{text},\mathbf{Emb}_{tx}\oplus\mathbf{Emb}_{clr})$. Note that, we only inject the disentangled embedding $\mathbf{Emb}_{tx}\oplus\mathbf{Emb}_{clr}$ into the first decoder layer to compute cross-attention maps, which shows better stylization performance as previous works proved wang2024instantstyle ; wang2024instantstyle_plus ; agarwal2023image_matte ; voynov2023ETI . Examples shown in Fig.˜6 (2nd cols.) demonstrate the effectiveness of our proposal.
Figure 5: We extract texture information and suppress gray expression through SVD reweighting, and color alignment is improved by RegWCT.‘Baseline’ denotes color-texture disentanglement with Image-Prompt Additivity only (i.e., without SVD and RegWCT). (a) shows the Roles of SVD and RegWCT within SADis. (b) Our approach enables more precise color alignment with $\mathcal{I}_{clr}$ via RegWCT, whereas direct WCT manipulation compresses texture details and high-frequency components. (c) RegWCT rectifies the latent distribution that is affected by WCT. The distribution of the image latent generated with RegWCT technique is closer to that of the image generated without WCT in the frequency domain, across diverse time steps.
@@ 3.3.3 Regularized Whitening-Coloring Transforms
WCT: whitening-coloring transformation. Solely applying our CTE method does not fully ensure a precise color palette match between the reference color image $\mathcal{I}_{clr}$ and the generated output (see Fig.5). We hypothesize that while CLIP embeddings capture high-level semantic information, they may not fully represent nuanced color distribution details. To better align the color distributions of the generated image with the color reference, we incorporate a whitening-coloring transform (WCT) li2017universal ; cho2019i2i_wct ; hossain2014wct applied to the noisy latent as $z_{t}^{wct}=\mathbf{WCT}(z_{t}^{{}^{\prime}})$ (detailed WCT formulas in the Supplementary). This approach allows for a more faithful color transfer by aligning the latent representations with the reference color palette, enhancing the stylistic consistency of generated images. Examples of applying the WCT transform during the T2I generation process are presented in Fig.˜5-(b), demonstrating an improved alignment of color palettes between the generated images and the reference color images. This enhancement visually confirms the effectiveness of WCT in achieving more accurate color transfer, resulting in a closer match to the desired color distribution.
@@
Noise Regularization for WCT. However, as shown in Fig.˜5-(b), the WCT process tends to compress texture details, distorting high-frequency information, which is also observed in previous works zhang2024real_bias ; li2023faster . This issue is attributed to the signal-leak bias everaert2024_signal_leak_bias and the inverted latent distribution gap from the inversion process zhang2024real_bias . These terms essentially describe the same underlying phenomenon: during training, the distribution of $z_{T}$—the latent representation after adding Gaussian noise to $z_{0}$—does not perfectly match a standard Gaussian. Instead, it always contains the leakage or some certain prior information towards the original image, causing the residual structure to persist even at high noise levels. This biased noise in the forward diffusion process leads to a slight misalignment in texture representation during generation. To address this, we propose adding a small-scale noise to the latent as $z_{t}=z_{t}^{wct}+\lambda\cdot\delta,\delta\sim\mathcal{N}(0,1)$, determined by a scale hyperparameter $\lambda$, to recover the lost high-frequency details. By introducing this noise regularization term, the regularized whitening-coloring transformation (RegWCT) achieves improved generative performance, better aligning both color and texture in the generated images with the reference images. Given a latent $z_{t}^{{}^{\prime}}$ at timestep $t$, the color rectified latent $z_{t}$ after RegWCT is formulated as $z_{t}=\left(1-\omega\right)z_{t}^{{}^{\prime}}+\omega\cdot\mathbf{RegWCT}(z_{t}^{{}^{\prime}})$, where $\omega$ acts as a balance weight. To prevent the inference trajectory from deviating too much from the original one, we only perform the RegWCT transformation during the intermediate steps as $T_{start},T_{end}]$. With all these techniques, including color-texture extraction (CTE) and Regularized WCT transformation (RegWCT), our method SADis achieves customizable and flexible Disentangled Stylized Image Generation (as shown in Fig.˜5 (a)-4th col.).
@@ 3.3.4 ControlNet-based real image stylization
Our method can be integrated with ControlNet zhang2023controlnet $\mathcal{CN}$ to facilitate content-based stylized image generation, as illustrated in Fig.˜1 (down) and Fig.˜7 (a). By using any pretrained ControlNet model (e.g. Canny-conditioned) as the base pipeline while maintaining all other hyperparameters consistent with SADis, we are able to significantly broaden the applicability of our method. More specifically, we have an input $\mathbf{I}_{c}$ as the content image, it is passed through the ControlNet $\mathcal{CN}(\mathbf{I}_{c})$ as conditions for T2I generation model as $z_{t-1}^{{}^{\prime}}=\mathcal{G}(z_{t},t,\mathcal{CN}(\mathbf{I}_{c}),\mathbf{Emb}_{tx}\oplus\mathbf{Emb}_{clr})$, where $z_{T}\sim\mathcal{N}(0,1)$ and the textual prompt as null $\mathcal{P}=\text{``''}$.
Figure 6: Disentangled Stylized Image Generation (DisIG) compared with baseline methods. For other approaches, we used GPT-4o to generate color descriptions based on the color reference image, incorporating these descriptions into the text prompts (details shown in the Supplementary). Our SADis accepts separate color and texture reference images to achieve flexible control.
@@ 4.1 Experimental Setups
Datasets. To ensure a fair comparison, we randomly select 40 images in total from the WikiArt artgan2018_wikiart and StyleDrop sohn2023styledrop datasets. For our method, SADis, each of these images serves as either a color reference or a texture reference to enable color-texture disentanglement. In contrast, since the comparison methods lack disentanglement capabilities, we treat each of these 40 images as a style reference for these methods, supplemented by auxiliary text prompts generated from GPT-4o achiam2023chatgpt4 as a strong captioning model. For quantitative comparisons, we used 20 images as the color reference set and 20 as the texture reference set. We also sampled 10 content prompts from StyleDrop, resulting in 4,000 stylized images per method to ensure fair and extensive comparisons with numerous images.
@@
Evaluation Metrics. We evaluate our method using multiple quantitative metrics. (1) For a general quality evaluation, we use CLIP-Score (CLIP) hessel2021clipscore to evaluate the T2I generation performance, specifically measuring the semantic consistency between the generated image and the text prompt. (2) For the evaluation of color attribute alignment, the MS-SWD he2024_MS_SWD metric specifically evaluates color distance between generated and reference images. Color histogram distance (C-Hist) computes the distance between both color histograms. The GPT-4o color score achiam2023chatgpt4 requests the GPT model to rate from 0-5 according to the color consistency between the color reference and the generated image. (3) For the texture quality, the Kernel-Inception Distance (KID) kid is assessing the quality of generated images by measuring the dissimilarity between the real and generated image distributions (We use feature dimension as 2048 and evaluate on the 4000 images). CLIP-I is used to evaluate the similarity between the texture image and the generated image in grayscale.
@@
Implementation Details. We build our method SADis upon the SDXL podell2023sdxl . To apply the image embeddings as additional conditions to the T2I model, we use the IP-Adapter ye2023ip-adapter pretrained projectors. Both of these models are based on the CLIP model radford2021clip , where the SDXL model utilizes the CLIP text encoder to generate textual embeddings and the IP-Adapter leverages the CLIP image encoder to extract image embeddings before the embedding projector. For the hyperparameters, we set $\lambda=0.01,T_{start}=0.8T,T_{end}=0.6T,\omega=0.5,\gamma=0.003,\beta=1.0$. For the time cost, we reported for 50-step DDIM. All the experiments are conducted on a single L40s GPU.
@@
Comparison Methods. We compare with several state-of-the-art tuning-free stylization approaches based on the T2I diffusion models: DEADiff qi2024deadiff , InstantStyle wang2024instantstyle , IP-Adapter ye2023ip-adapter , CSGO xing2024csgo , DreamStyler ahn2023dreamstyler , StyleDrop sohn2023styledrop and Artist jiang2024artist . For these baseline methods, we utilize the GPT-4o achiam2023chatgpt4 model to generate captions from the color reference image, which are appended to the original textual prompt to guide the T2I generation. Additionally, SDXL podell2023sdxl and Artist model are included in comparisons by generating captions from both the color and texture reference images, appending these captions to the textual prompts to guide the generation process. We include much more experimental setup details, qualitative and quantitative results in the Supplmentary Material.
@@ 4.2 Experimental Results
Qualitative Comparison. In our comparison of generation quality, visual results under the Disentangled Stylized Image Generation (DisIG) scenario are shown in Fig.˜6 and  Fig.˜7. While baseline methods manage to capture some style information from the style reference image (the texture image for SADis), they fail to represent the color palette provided by the color reference accurately. The IP-Adapter and InstantStyle methods even retain original layouts and elements from the style images, deviating from the intended stylization. Other baselines, such as Artist jiang2024artist and DEADiff qi2024deadiff , exhibit poor generation quality, reflecting a limited understanding of both color and texture attributes. In contrast, our method, SADis, allows for both texture and color reference inputs, offering precise, controllable T2I stylization. The generated images by SADis exhibit significantly improved color fidelity and texture detail compared to other approaches, demonstrating its effectiveness in disentangled style attribute transfer. To further validate SADis’s robustness, we conduct two tests: (1) swapping the color and texture reference images as shown in Fig.˜7 (c), and (2) modifying their saturation and illumination levels, as shown in Fig.˜7 (d). Unlike existing approaches such as InstantStyle, which cannot preserve image content under these conditions (detailed in Supplementary), SADis successfully applies color and texture independently while keeping the content largely unchanged.
Table 1: Quantitative Comparison with existing image stylization methods. The best and second-best numbers are marked with bold and underlined respectively.
Method CLIP Color Texture Time Cost (s) User study (%)
@@
SDXL podell2023sdxl 0.272 9.51 1.20 3.01 0.69 0.08 9.26 16.99 5.84 11.65
IPAdapter ye2023ip-adapter 0.233 11.54 1.23 2.84 0.84 0.043 9.52  6.08 22.54 14.55
InstantStyle wang2024instantstyle 0.261 12.53 1.32 2.80 0.74 0.056 9.31  4.81 24.94 13.10
Artist jiang2024artist 0.269 10.48 1.39 2.82 0.69 0.089 12.32  9.38 1.79 3.40
DEADiff qi2024deadiff 0.267 11.20 1.24 2.73 0.69 0.087 1.86  4.57 2.59 2.67
StyleDrop sohn2023styledrop 0.275 13.52 1.43 2.67 0.70 0.054 6.91  5.07 3.57 5.34
DreamStyler ahn2023dreamstyler 0.277 12.17 1.26 2.39 0.71 0.060 5.23  4.67 3.57 5.39
CSGO xing2024csgo 0.280 14.25 1.36 2.63 0.69 0.071 15.99  6.59 9.73 6.06
SADis (Ours)
0.281 5.57 0.96 3.34 0.74 0.049 10.30 41.83 25.42 37.84
Table 2: Ablation by removing each components in our method SADis.
Method Color Texture Time cost (s)
@@
[…]
SADis (Ours)
7.27 0.96 0.74 $\approx 10.30$
[…]
@@
Quantitative Comparison. A detailed quantitative comparison is provided in Tab.˜1 to further support our findings. Our method, SADis, retains text-image alignment quality at a level comparable to the base SDXL model podell2023sdxl , as indicated by the CLIP score. This comparison demonstrates that SADis preserves alignment with the textual prompt better than other methods. Regarding color alignment, SADis significantly outperforms other approaches. For texture representation, the IP-Adapter-based methods, including InstantStyle wang2024instantstyle and SADis, show superior performance over other baselines. They achieve higher texture-related metrics at the expense of color fidelity and text-image alignment quality. The high texture scores for IP-Adapter can be explained by the fact that they control all cross-attention layers; However, this leads to high-semantic content leakage (like the same stars in the Van Gogh bear in Fig.˜6). We decided to prevent this by only controlling the cross-attention of the low-level layers following InstantStyle wang2024instantstyle ; agarwal2023image_matte ; voynov2023ETI ; Therefore, our texture results are close to those of InstantStyle, but at the same time we greatly improve color fidelity. In addition, their generated images follow the same structure as the original texture image, showing that they do not disentangle structure from texture but overfit to the contents of the texture images.
@@
Ablation Study. The ablation study for each component of SADis is listed in Tab.˜2. Results show that SVD rescaling in the CTE process and the noise-regularized RegWCT techniques significantly enhance color alignment while only slightly reducing texture precision, which is almost undetectable in the CLIP-I metric. The trade-off between texture and color alignment is optional for users, allowing them to adjust the balance based on the requirements.
@@
User Study. To better assess alignment with human preferences, as shown in Tab.˜1(right), we conducted a user study with 24 participants (30 sextuplets/user), collecting 720 data for each method. Each was asked to “select the best image from pairs of generated images, taking into account overall quality (considering both color and texture alignments), color alignment, and texture alignment, respectively.” Our method outperformed other baseline approaches with at least a 20% improvement. This demonstrates the potency of SADis and its high alignment with human preference.
4.3 Additional Features